# Novel Application of Neutrinos to Evaluate U.S. Nuclear Weapons Performance

### v1

@@ III.0.2 Prompt Fission Yield Assumption
On the relatively slow timescale for neutrino emission from an nuclear explosion, the neutrinos are produced by the equivalent of a neutron pulse inducing the fission, and the shape, magnitude and timing of the spectrum is determined by the direct population and subsequent beta-decay of the fission fragments. The fission products (FP) are produced both directly in fission, with an independent yield (), and from the beta-decays of the more neutron rich fission products of the same mass. In order to extract the time dependent energy spectrum of antineutrinos emitted through the beta decays of the prompt FPs, we make the assumption that the FP’s were populated and distributed equally to the known FP independent yields for a given fuel isotope by an approximately zero-duration event at what we define to be . We then track the decays of each FP along each possible decay chain until a distribution of quasi-stable (i.e. sufficiently long lived) isotopes and isomers are populated from the initial distribution. We obtained each species’ half-life () and decay branching ratios () for each decay from the RIPL-3 nuclear reaction database [17]. We determine the time-dependent yield of each species (denoted as simply ) by solving a coupled, linear, initial value problem for which the instantaneous change in the time dependent yield of an isotope with atomic number , mass , and isomer index is given by
@@
Here, is the decay rate of the species which has halflife . The sum over represents the total contribution of all possible species which decay into the LHS species , and the factor represents the branching ratio for which the parent species undergoes the specific decay to the LHS species. The sum includes such processes as transitions between isomer states, beta decay, and beta decay with emitted neutrons.
@@
The prompt fission assumption enters the model through the initial condition (i.e. at the yields are given by the independent yields of the fissioning isotope) and the absence of a source term in eq. 3. If the fissioning material changes composition or fissioning rate on a timescale which is comparable to the typical decay rates present in eq. 3, then a source term would be included in the evolution of the yields.
The instantaneous total antineutrino energy spectrum is then the weighted sum over all of the beta decay spectra from the present decaying nuclei 222We have further assumed that all considered beta decays are allowed. A more detailed analysis including shape factor corrections to forbidden decays will be presented in a future work., and the weights are a product of the time dependent yield fractions determined above, the total beta decay branching ratio for the decay of each species, and each species decay rate. The total number of antineutrinos emitted () per unit time is the time derivative of the energy-integrated spectrum. We plot the total number of emitted antineutrinos per second per fission in the left panel of Fig. 3. Because the decay rate is larger for higher energy antineutrinos, those emitted at earlier times have a correspondingly higher average energy, and we plot the instantaneous average antineutrino energy as the fission yield distribution evolves in the right panel of Fig. 3.
@@
A reactor operating at steady state fissions actinide isotopes at a nearly constant rate. Once the distribution of isotopes in the reactor reaches its equilibrium state, the rate at which any individual isotope decays is equal to the rate at which it is populated both directly from fissions and from the decay of other isotopes. For unstable isotopes the value to which a yield saturates is proportional to the ratio of the steady state fissioning rate to the specie’ half-life and the species cumulative yield. The cumulative yield is simply the sum of the independent yields of all of the iosotopes in the decay chain “above” the isotope under consideration (suitably weighted by the branching ratios along the decay chain). The cumulative yields for fast and 14 MeV neutron induced fissioning of 239Pu are known [19] and are utilized to model neutrino fluxes from reactors operating under equilibrium conditions [13]. Under such steady state conditions, the effective cross sections, number of neutrinos per fission, and average neutrino energy all become time independent and we provide them for ease of comparison in Tables 1, 2, and 3. The expected number of antineutrinos from a given reactor with a known fuel composition can be calculated from the cumulative yields of the reactor fuel components and is found to be for a typical Uranium power generation reactor [13]. The number of antineutrinos per fission we find for the prompt fission case is substantially lower at (see Table 1), however the spectrally weighted cross section () is of comparable magnitude: for 239Pu fissioning in a reactor environment the Daya Bay Collaboration measured cm2 per fission [20], to be compared with Table 2. In a reactor, the majority of the antineutrinos emitted are of energies below the IBD threshold and so do not produce a signal in an IBD detector resulting in an overall comparable detector response for both the reactor and nuclear explosion cases.
@@ III.0.3 Systematic Errors in Nuclear Explosive Yield
The neutrino measurement error depends on the number of neutrino events detected (statistical), and systematic errors coming from detection threshold energy resolution and reconstruction efficiency (systematic). For statistics in the thousands of events and reasonable assumptions on systematic errors, we can expect a total error on the measured neutrino neutrino flux of about 4%, or less.
A more uncertain error comes from extrapolating the measured neutrino flux back to a test device explosive yield. A few of the assumptions were discussed above. These include errors from the prompt fission yield assumption, the detailed evolution of the neutron flux that is inducing fission, uncertainties in evaluated nuclear data and reaction libraries, and fission product database uncertainties (ENDF [19, 17]). These assumptions need to be tested and measured in a control experimental setup as described in the next section.
@@ IV Near-term Pulsed Reactor Demonstrator Experiment
In the absence of nuclear weapons testing for the foreseeable future, many of the technical and theoretical challenges of a full nuclear test could be mitigated in the next few years with a smaller scale moderate cost demonstration detector. The plan is to deploy a 20 tons fiducial mass (25 tons total) IBD detector placed near a pulsed reactor, such as the Texas A&M TRIGA 1GW-10 millisecond pulsed facility. The short duty cycle and repeatability of pulses would provide critical real condition testing and knowledge which would be valuable for planning a possible real test shot in the future. Furthermore, the measured neutrino rate per pulse as a function of time would provide unique constraints on fission databases, equilibrium assumptions, and yield errors. Such a near-term demonstrator experiment could be funded quickly and at a reasonable cost.
Typically, a pulsed reactor can generate fissions in a single pulse, thus many pulses are needed to achieve the number of fissions required for the demonstration. By recording events triggered in a 1000 seconds timing window following a reactor pulse, backgrounds can be further reduced and a decay curve of the fission products can be measurement. The close proximity of the detector and the ability offline to superimpose multiple pulses on the same pulse trigger allows an accurate simulation of the event rate pile-up in a nuclear explosion. In Section IV.0.1 and IV.0.2, we will discuss various pulsed neutron sources that we have explored and number of neutrinos that we expect to see at each facility. Section IV.0.3 details the technical measurements to be made and Section IV.0.4 will describe our preliminary detector design. The size of the detector has to allow for sufficient statistics based on the total number of fissions that we expect at each facility.
@@ IV.0.1 Pulsed Neutrino Sources
One of the pulsed reactor sources that we have identified are TRIGA reactors, which can run in steady-state and pulsed mode [21]. The TRIGA reactors can produce a brief supercritical pulse and generate a tremendous amount of neutrino flux over 10’s of seconds, which is ideal for separation from the fast neutron and gamma related backgrounds and random backgrounds. The Texas A&M (TAMU) TRIGA reactor has been identified as a potential neutrino source. The 20% 235U enriched reactor has a steady-state fission rate of 1-megawatt and can produce a peak power 250-megawatt pulse with a width of up to 40 ms, producing fissions per pulse [21]. The reactor also has large floor space and experimental facilities to deploy a moderate size detector as close as 4 m, including shielding, from the reactor core (see Figure 4). The pulsed reactor at Sandia National Laboratory [22] has even higher pulsed power characteristics as TAMU, greater than fissions per pulse. We will also be investigating the logistical feasibility of mounting a neutrino detector at the Sandia pulsed reactor.
@@ IV.0.2 IBD and Background event rates at TAMU TRIGA
The expected number of fissions per pulse at the TAMU TRIGA during normal operations is fissions/pulse. Due to small neutrino rates per pulse, analysis will only take place between 0.1 and 100 seconds after a pulse to reduce steady state backgrounds (the DAQ window will still be active for at least 1000 seconds). Calculating for a prompt fissioning of 235U from thermal neutrons in this time range gives cm2. Given a detector distance of 4 m (center of reactor to center of detector), and the same number of free per detector ton from Section II, then the rate is 0.0066 neutrinos/ton/pulse (this already includes threshold efficiencies). For 2000 pulses an operation year [23], a 20-ton fiducial detector discussed below, and 80% detection efficiency (Section II), then the IBD event rate is 210 events/year. This is a small but reasonable amount of events to test the various technical challenges detecting neutrinos from a short burst. A three year run will provide a 4% statistical error on the measured pulsed power yield. Given the higher power, the event rate at Sandia could be much larger if the detector can be placed at a similar nearby distance as at TAMU, and is currently being investigated.
@@
Shielding will be added to reduce both fast and thermal neutrons and gamma-rays from the reactor pulse. However, it is still anticipated that a significant rate will make it to the detector at such a close distance. The strategy to mitigate this prompt component is to cut out from analysis the first 100 millseconds of data. This will result in only a 2% reduction in neutrino signal since the majority come from slower fission decays (see Table 2).
@@
Accidental backgrounds from cosmic rays (500 Hz) and PMT beta-gamma’s (2 kHz) are easily rejected by veto and large charge cuts and will create an analysis dead-time of only 0.25%. Steady state background from activation and other external source of MeV gamma-rays is expected to be about 10 kHz (based on the 10-ton LAr CCM detector measurements at Lujan neutron source [24]). Adding appropriate shielding including borated poly (for neutron capture) and steel (for gamma absorption) should reduce this to 3.5 kHz, again from CCM experience. After 1.8 MeV energy threshold and fiducial cuts this can be reduced to 100 Hz. Finally, accidental neutrons from cosmogenic and PMT glass sources need to be considered, but are expected to be small. At this level the accidental background rates to IBD (for a 10 micro-second neutron capture time) from external gamma-ray backgrounds is 0.1%. Using a 100 second DAQ window pulsed 2000 times is 200 events in a one year run. As this is steady state it can be measured precisely by strobe and pre-pulse data and subtracted from signal and represents a 10% subtraction error on the expected signal. However, adding more sheilding and improved analysis to control the rate is important to reducing the statistical error in the subtraction.
@@ IV.0.3 IBD Technical Tests Measurements at TAMU TRIGA
The primary purpose of running a IBD detector at a pulsed reactor is to test and understand the characteristics and issues related to IBD neutrino detection from a short pulsed source. In the past all IBD neutrino detector experiments have been performed at steady state reactors. The capability of these detectors have been well characterized and proven. However, a pulsed source such as an nuclear weapons test has some unique challenges. The following is a list of unique tests to be performed for a IBD detector at a pulsed reactor source,
Most of the neutrinos arrive within 1000 seconds. The preferred run mode is an unbiased trigger (collecting all digitization data for all PMT’s over the entire 1000 seconds) puts an extremely large data throughput load on the DAQ (1000 TBytes/pulse). The heavy data throughput will require new solutions such as faster readout hardware, more DAQ computers, and increased fast data storage buffer that all need to be tested and implemented. If an unbiased trigger cannot be implemented, then a event trigger based on detector activity will have to be studied and implemented. There can be a loss of physics if this is not done carefully and tested before a potential one-time shot.
Large pile up of IBD events from a weapons test up to a 1000 second interval will present reconstruction challenges. An IBD event is identified by a prompt position then followed by a capture neutron on Gd within about 10 micro-seconds. With multiple events happening together (up to many thousands from a nuclear explosion), uniquely tagging a positron with a capture neutron becomes more difficult. At a low rate steady state reactor this is not an issue, but for an intense pulsed source it becomes critical. Any one pulse from the TRIGA reactor will have minimal overlap, but after collecting a few years worth of many pulses, these events can be overlapped to produce a data stream that mimics the intensity with a bomb. Analysis strategies can then be developed and tested before a one-time shot. A potential outcome of this study is the realization that capturing the IBD neutron with Gd will only confuse the event reconstruction and is best not added to the scintillator.
@@
Starting with the IBD neutrino rate measurement, we will apply reactor fission codes (such as ENDF [19]) to extrapolate back to the reactor power per pulse, which is known to 1% level [23]. This is important for testing the various methods and codes necessary for bomb test yield predictions using neutrino rates. From steady state reactor measurements in the past we know this error to 3-4%. The TAMU TRIGA reactor runs half the time in steady state mode which will allow a measurement of the steady state neutrino rate (this is estimated to be over 10,000 events/year, albeit with much higher background rate). Comparison of the steady state rate to pulsed mode with the same reactor and detector could yield useful information on equilibrium fission rates versus pulsed dynamics.

### v2

@@ III.0.2 Prompt Fission Yield Assumption
On the relatively slow timescale for neutrino emission from an nuclear explosion, the neutrinos are produced by the equivalent of a neutron pulse inducing the fission, and the shape, magnitude and timing of the spectrum is determined by the direct population and subsequent -decay of the fission fragments. The fission products (FP) are produced both directly in fission, with an independent yield (denoted ), and from the -decays of the more neutron rich fission products of the same mass. In order to extract the time dependent energy spectrum of antineutrinos emitted through the -decays of the prompt FP’s, we make the assumption that the FP’s were populated and distributed equally to the known FP independent yields for a given fuel isotope by an approximately zero-duration event at what we define to be . We then track the decays of each FP along each possible decay chain until a distribution of quasi-stable (i.e. sufficiently long lived) isotopes and isomers are populated from the initial distribution. We obtained each species’ half-life () and decay branching ratios () for each decay from the RIPL-3 nuclear reaction database [17]. We determine the time-dependent yield of each species (denoted as simply ) by solving a coupled, linear, initial value problem for which the instantaneous change in the time dependent yield of an isotope with atomic number , mass , and isomer index is given by
@@
Here, is the decay rate of the species which has half-life . The sum over represents the total contribution of all possible species which decay into the left hand side (LHS) species , and the factor represents the branching ratio for which the parent species undergoes the specific decay to the LHS species. The sum includes branching over all possible transitions between species such as: internal conversions between isomer states, beta decay, and beta decay with emitted neutrons.
@@
The prompt fission assumption enters the model through the initial condition (i.e. at the yields are given by the independent yields of the fissioning isotope) and the choice in eq. 4. If fissioning is extended in time, then the source term must be accounted for in tracking the evolution of each species’ time dependent yield.
@@
A fission nuclear reactor operating at steady state fissions actinide isotopes at a nearly constant rate, resulting in taking a constant value. After the reactor reaches steady state, the rate at which any individual isotope decays is equal to the rate at which it is populated both directly from fissions and from the decay of other isotopes. For unstable isotopes the value to which a yield saturates is proportional to the ratio of the steady state fissioning rate to the species’ half-life and the species cumulative yield. The cumulative yield is simply the sum of the independent yields of all of the isotopes in the decay chain “above” the isotope under consideration (suitably weighted by the branching ratios along the decay chain). The cumulative yields 239Pu and 235U are known [18] and are utilized to model neutrino fluxes from reactors operating under equilibrium conditions [19]. Under such steady state conditions, the dynamic cross sections, number of neutrinos per fission, and average neutrino energy all become time independent and we provide them for ease of comparison in the rightmost columns labeled “S.S.” in tables 1, 2, and 3. In the evaluation of the quantities we report, we include in the antineutrino emission processes only the -decays of nuclei; we have not included the decay of any present unbound neutrons in the S.S. system.
Once the time-dependent fission product yields are known, the instantaneous total antineutrino energy spectrum (denoted ) is then the weighted sum over all of the -decay spectra from the present decaying nuclei 222We have further assumed that all considered beta decays are allowed. A more detailed analysis including shape factor corrections to forbidden decays will be presented in a future work., and the weights are a product of the time dependent yield fractions determined above and the total -decay branching ratio for the decay of each species. If each species’ antineutrino energy spectrum is denoted then the total instantaneous spectrum is
The instantaneous average neutrino energy is then
The total number of antineutrinos emitted () per unit time (e.g. the total antineutrino emission rate) is the sum of the -decay rates of the individual species such that
Here is the same as in eq. 4, but is the branching ratio with which the species specifically -decays. We plot the total number of emitted antineutrinos per second per fission in the left panel of Fig. 3. Because the decay rate is larger for higher energy antineutrinos, those emitted at earlier times have a correspondingly higher average energy, and we plot the instantaneous average antineutrino energy as the fission yield distribution evolves in the right panel of Fig. 3.
If we consider some window of time defined as an interval , then we let number of antineutrinos emitted by -decaying species be denoted as and the total number of fissions undergone by a fissioning actinide in that window is denoted as . Then the ratio is just given as
While a somewhat simple restatement, casting the ratio of antineutrinos-per-fission in this form allows us to more clearly elucidate the difference in this ratio for pulsed vs steady-state events. When the number of fissions varies in the time window, one generally must know in detail both how the fission product yields vary as a function of time and how the fissions vary in time. For a fission pulse at , and the neutrino emission rate will have a nontrivial time dependence as the fission product yields start their evolution at their independent yield values at and evolve toward a stable final distribution at (assuming that is chosen to be sufficiently far into the future). Thus, generally the total number of antineutrinos per fission that one expects to be emitted from some fission environment is a nontrivial function of that environment’s initial condition, subsequent evolution, and the specific window of time over which one considers.
The expected number of antineutrinos from a given reactor with a known fuel composition can be calculated from the cumulative yields of the reactor fuel components and is found to be for a typical Uranium power generation reactor. The number of antineutrinos per fission emitted in the first seconds after the prompt fission case we find to be substantially lower (see Table 1), however the spectrally weighted, and time integrated, cross section () is of comparable magnitude: for 239Pu fissioning in a reactor environment the Daya Bay Collaboration measured cm2 per fission [21], while for 235U they measured cm2 per fission to be compared with Table 2. A reactor fissioning in steady-state emits antineutrinos at an approximately constant rate. The majority of these in any arbitrary window of time are of energies below the IBD threshold and so do not produce a signal in an IBD detector. For the pulsed case, most of the neutrinos below the IBD threshold are emitted at late time, and this results in the observation that the spectrally averaged and time integrated cross sections we present in Table 2 are of a comparable values to those of a steady state reactor.
@@ III.0.3 Systematic Errors in Nuclear Explosive Yield
The neutrino measurement error depends on the number of antineutrino events detected (statistical), and systematic errors coming from detection threshold energy resolution and reconstruction efficiency (systematic). For statistics in the thousands of events and reasonable assumptions on systematic errors, we can expect a total error on the measured antineutrino flux of about 4%, or less.
A more uncertain error comes from extrapolating the measured antineutrino flux back to a test device explosive yield. A few of the assumptions were discussed above. These include errors from the prompt fission yield assumption, the detailed evolution of the neutron flux that is inducing fission, uncertainties in evaluated nuclear data and reaction libraries, and fission product database uncertainties (ENDF [18, 17]). These assumptions need to be tested and measured in a controlled experimental setup as described in the next section.
@@ IV Near-term Pulsed Reactor Demonstrator Experiment
In the absence of nuclear weapons testing for the foreseeable future, many of the technical and theoretical challenges of a full nuclear test could be mitigated in the next few years with a smaller scale moderate cost demonstration detector. The plan is to deploy a 20 ton fiducial mass (25 tons total) IBD detector placed near a pulsed reactor, such as the Texas A&M TRIGA 1GW-10 millisecond pulsed facility. The short duty cycle and repeatability of pulses would provide critical real condition testing and knowledge which would be valuable for planning a possible real test shot in the future. Furthermore, the measured neutrino rate per pulse as a function of time would provide unique constraints on fission databases, equilibrium assumptions, and yield errors. Such a near-term demonstrator experiment could be prepared quickly and at a reasonable cost.
Typically, a pulsed reactor can generate fissions in a single pulse, thus many pulses are needed to achieve the number of fissions required for the demonstration. By recording events triggered in a 1000 seconds timing window following a reactor pulse, backgrounds can be further reduced and a decay curve of the fission products can be measured. The close proximity of the detector and the ability offline to superimpose multiple pulses on the same pulse trigger allows an accurate simulation of the event rate pile-up in a nuclear explosion. In Section IV.0.1 and IV.0.2, we will discuss various pulsed neutrino sources that we have explored and the number of antineutrinos that we expect to see at each facility. Section IV.0.3 details the technical measurements to be made, Section IV.0.4 will describe our preliminary detector design, and IV.0.5 discusses detector response and simulations. The size of the detector has to allow for sufficient statistics based on the total number of fissions that we expect at each facility.
@@ IV.0.1 Pulsed Neutrino Sources
One of the pulsed reactor sources that we have identified is TRIGA reactors, which can run in steady-state and pulsed mode [22]. The pulsed mode is accomplished by rapidly firing the control rod out of the core within a fraction of a second. The reactor quickly produces a super-criticality pulse from a low power baseline. As the temperature of the fuel increases, its reactivity decreases and allows the reactor to return to normal levels, the reactor is then automatically scrammed after a few seconds to clip the power tail for operational convenience [23]. This brief supercritical pulse of fission can generate a tremendous antineutrino flux over 10’s of seconds, which is ideal for separation from the prompt fast neutron and gamma related backgrounds and steady state random backgrounds.
The Texas A&M (TAMU) TRIGA reactor has been identified as a potential neutrino source. The 20% 235U enriched reactor has a steady-state fission rate of 1-megawatt and can produce a peak power 250-megawatt pulse with a width of up to 40 ms, producing fissions per pulse [22]. The reactor also has large floor space and experimental facilities to deploy a moderate size detector as close as 4 m, including shielding, from the reactor core (see Figure 4). The pulsed reactor at Sandia National Laboratory [24] has even higher pulsed power characteristics as TAMU, greater than fissions per pulse. We will also be investigating the logistical feasibility of mounting a neutrino detector at the Sandia pulsed reactor.
@@ IV.0.2 IBD and Background event rates at TAMU TRIGA
The expected number of fissions per pulse at the TAMU TRIGA during normal operations is fissions/pulse. Due to small neutrino rates per pulse, analysis will only take place between 0.1 and 100 seconds after a pulse to reduce prompt and steady state backgrounds (the DAQ window will still be active for at least 1000 seconds). Calculating , from Table 2, for a prompt fissioning of 235U from thermal neutrons in this time range gives cm2. Given a detector distance of 4 m (center of reactor to center of detector), and the same number of free per detector ton from Section II, then the rate is 0.0068 antineutrinos/ton/pulse (this already includes threshold efficiencies). For 2000 pulses per operational year [25], a 20-ton fiducial detector discussed below, and 80% detection efficiency (Section II), then the IBD event rate is 218 events/year. This is a small but reasonable amount of events to test the various technical challenges of detecting neutrinos from a short burst. A three year run will provide a 4% statistical error on the measured pulsed power yield. Given the higher power, the event rate at Sandia could be much larger if the detector can be placed at a similar nearby distance as at TAMU.
@@
Shielding will be added to reduce both fast and thermal neutrons and gamma-rays from the reactor pulse. However, it is still anticipated that a significant number will make it to the detector at such a close distance. The strategy to mitigate this prompt component is to cut out from analysis the first 100 milliseconds of data. This will result in only a 2.7% reduction in antineutrino signal since the majority come from slower fission decays (see Table 2).
@@
Accidental backgrounds from cosmic rays (500 Hz) and PMT beta-gamma’s (2 kHz) are easily rejected by external muon paddle detectors and PMT large charge cuts and will create an analysis dead-time of only 0.25%. Steady state background from activation and other external source of MeV gamma-rays is expected to be at most 10 kHz (based on the 10-ton LAr CCM detector measurements at Lujan neutron source [26]). Adding appropriate shielding including borated poly (for neutron capture) and steel (for gamma absorption) should reduce this to 3.5 kHz, again from CCM experience. After the 1.8 MeV energy threshold and fiducial cuts this can be reduced to 100 Hz. Finally, accidental neutrons from cosmogenic sources [27] (mimic IBD via neutron capture coincidence with random gamma-ray or neutron scatter producing gamma-ray then thermalizing and capture) and PMT glass [28] sources are estimated to be Hz rate. At this level the accidental background rates to IBD (for a 10 microsecond neutron capture time) from external backgrounds is 0.1%. Using a 100 second DAQ window pulsed 2000 times gives 200 events in a one year run. As this is steady state it can be measured precisely by strobe and pre-pulse data and subtracted from signal and represents a event statistical subtraction error, or about on the expected signal. However, adding more shielding and improved analysis to control the rate is important to reducing the statistical error in the subtraction.
@@ IV.0.3 IBD Technical Tests Measurements at TAMU TRIGA
The primary purpose of running an IBD detector at a pulsed reactor is to test and understand the characteristics and issues related to IBD neutrino detection from a short pulsed source. In the past all IBD neutrino detector experiments have been performed at steady state reactors. The capability of these detectors have been well characterized and proven. However, a pulsed source such as an nuclear weapons test has some unique challenges. The following is a list of tests to be performed for an IBD detector at a pulsed reactor source,
Most of the antineutrinos arrive within 1000 seconds. The preferred run mode is an unbiased trigger (collecting all digitization data for all PMT’s over the entire 1000 seconds) which puts an extremely large data throughput load on the DAQ (1000 TBytes/pulse). The heavy data throughput will require new solutions such as faster readout hardware, more DAQ computers, and an increased fast data storage buffer that all need to be tested and implemented. If an unbiased trigger cannot be implemented, then an event trigger based on detector activity will have to be studied and implemented. There can be a loss of physics if this is not done carefully and tested before a potential one-time shot.
Large pile up of IBD events from a weapons test up to a 1000 second interval will present reconstruction challenges. An IBD event is identified by a prompt positron then followed by a capture neutron on Gd within about 10 microseconds. With multiple events happening together (up to many thousands from a nuclear explosion), uniquely tagging a positron with a correlated capture neutron becomes more difficult. At a low rate steady state reactor this is not an issue, but for an intense pulsed source it becomes critical. Any one pulse from the TRIGA reactor will have minimal event overlap, but after collecting a few years worth of many pulses, these events can be overlaid to produce a data stream that mimics the higher intensity with a bomb. Analysis strategies can then be developed and tested before a one-time shot. A potential outcome of this study is the realization that capturing the IBD neutron with Gd will only confuse the event reconstruction and is best not added to the scintillator.
@@
Starting with the IBD antineutrino rate measurement, we will apply reactor fission codes (such as ENDF [18]) to extrapolate back to the reactor power per pulse, which is known to 1% level [25]. This is important for testing the various methods and codes necessary for bomb test yield predictions using antineutrino rates. From steady state reactor measurements in the past we know this error to 3-4%. The TAMU TRIGA reactor runs half the time in steady state mode which will allow a measurement of the steady state antineutrino rate (this is estimated to be over 10,000 events/year, albeit with much higher background rate). Comparison of the steady state rate to pulsed mode with the same reactor and detector could yield useful information on equilibrium fission rates versus pulsed dynamics.